OpenStudy (anonymous):

4x-2y=-10 and -6x+3y=12

OpenStudy (anonymous):

which method do u want to use?

OpenStudy (anonymous):

its a question im trying to figure out . Graphing 2 x 2 .

OpenStudy (anonymous):

you want to get the value for x and y that satisfy the 2 eqs

OpenStudy (anonymous):

So for example , if the quesion says . Select all the terms that describe the following system: and it listed a One solution b Coinciding c Intersecting d Inconsistent e No solution f Independent g Dependent h Parallel. Do i still figure the equation out ?

OpenStudy (anonymous):

wanna do it together or u want the answer?

OpenStudy (anonymous):

Can you just give me the answer.

OpenStudy (anonymous):

e No solution (4x-2y=-10 )*3=12x-6y=-30 (-6x+3y=12)*2=-12x+6y=24 add both equations 0 + 0 = 6 0 = 6 no solution

OpenStudy (anonymous):

Thank You !

OpenStudy (anonymous):

no problem; its better for u to understand how to do it

OpenStudy (anonymous):

Yes . it is.